# Unpacking Personal(?!) Health Informatics: An Investigation of Awareness, Understanding, & Leveraged Utility in India

### v1

@@ Abstract.
Personal Health Informatics (PHI), which leverages digital tools and information systems to support health assessment and self-care, holds promise for empowering individuals and transforming healthcare delivery. However, barriers to its adoption remain underexplored in the Indian context. This study investigates PHI adoption among Indian users and stakeholders using a multi-method approach. An awareness survey (n = 87) examined the usage of wearables and general PHI engagement, followed by semi-structured interviews (n = 22) that explored motivations, usage patterns, and health information sources. Qualitative analysis revealed that while PHI is valued for health monitoring and shared/collective care, its adoption is hindered by factors such as low health literacy, usability challenges, and mistrust in digital health platforms. Further stakeholder interviews and co-design workshops informed the development of a Figma-based prototype, which was evaluated for usability. Based on these findings, we offer design recommendations for an integrated, user-controlled PHI platform featuring accessible analytics and verifiable health information. Our insights highlight the socio-technical challenges of PHI adoption in India and underscore the need for reliable, user-centric solutions to support proactive healthcare.
@@ 1. Introduction
Personal health informatics (PHI) involves using information systems and tools to assess health and wellbeing and holds significant promise for empowering individuals and transforming healthcare by enabling active self-management. Globally, PHI leverages technologies—from smartphone applications to wearable devices such as smartwatches and fitness trackers—facilitating continuous health monitoring (Shajari et al., 2023). Factors like high health literacy, technology familiarity, and a proactive healthcare ethos drive PHI adoption in the Western context (Kyytsönen et al., 2023; Chandrasekaran et al., 2020; Zhang and Shahriar, 2020; Huang et al., 2022; Zhan et al., 2024; Zaher et al., 2024). Emerging evidence (Chmielewski and Meyer, 2024) in the Western context suggests that healthcare providers and users are beginning to view PHI as a medium for a more engaged care—seeking personal insights and opportunities to share data with healthcare providers for more proactive management.
@@
In India, however, infrastructural, sociocultural, and technological challenges intersect and constrain PHI’s potential. Digital health tools, including wearables, are often perceived as fashion accessories or status symbols rather than integrated components of comprehensive health management (Misra et al., 2023). Moreover, cultural practices, family dynamics, and skepticism towards advanced technologies, such as AI-driven solutions, further influence PHI usage and adoption (Hoque and Bao, 2015; Zhang et al., 2022; Creaser et al., 2022; Patil et al., 2022; Reis et al., 2024; Kapania et al., 2022; Wisetsri and Vijai, 2021). Given the cultural dynamics of the Indian context, a significant gap in its adoption and use can be filled by examining and understanding how PHI is currently integrated and utilized in daily health practices in India and shared with family and loved ones.
To address these challenges, we investigated the adoption, integration, and perception of PHI among Indian users and stakeholders using a multi-method approach. We conducted a survey (n=87) to assess awareness and usage of smart wearables, followed by interviews (n=22) exploring user motivations, usage patterns, and health information sources. Additional interviews with key stakeholder pairs—couples, intergenerational family groups, and healthcare providers—and co-design workshops (n=12) employing card sorting examined shared care practices and data-sharing preferences. Further, we evaluated a figma prototype to understand its usability and get design recommendations from the participants on what, how, and when they would like to share their health information/analytics with different stakeholders.
@@
Guided by Murnane et al.’s ecological model of PHI (Murnane et al., 2018) and Bronfenbrenner’s Ecological Systems Theory (Bronfenbrenner, 1974), we investigate both intrinsic and extrinsic factors influencing PHI adoption in India. Specifically, we ask: (1) How do Indians adopt, integrate, and perceive PHI in their health management? (2) What dictates PHI’s effective utilization and perceived utility, and how can we improve them?
@@
Our findings offer a deeper understanding of PHI challenges and opportunities in India and provide a Proof-of-Concept evaluation to strengthen our recommendations further. Through this work, we lay the groundwork for a more reliable and integrated PHI system supporting personal and shared care management, culturally appropriate to the Indian context and the global south.
@@ 2. Related Work
The body of work on personal health informatics within the broader HCI community, though established, has seen significant expansion in recent years, driven by the rising prevalence of chronic diseases associated with sedentary lifestyles and advancements in digital health technologies (Anderson and Durstine, 2019). These factors have propelled research toward innovative health management solutions and broadened the adoption of PHI systems. The rapid increase in chronic health conditions and an expanding digital landscape make personal health informatics especially critical in India. This work aims to contextualize the evolution of PHI within the Indian healthcare landscape, addressing its unique challenges and opportunities. This section examines the existing scholarship through two critical lenses: (1) personal health informatics in HCI and digital health tracking and (2) the barriers and enablers of PHI in the Indian context. These lenses offer complementary perspectives, highlighting technological and user-centric design considerations while incorporating them into India’s sociocultural health practices.
2.1. Personal Health Informatics in HCI
@@
Personal health informatics (PHI) within HCI has evolved from individual tracking (Rooksby et al., 2014) to encompassing broader ecological models (Epstein et al., 2015) that account for social, structural, and cultural factors. Murnane et al.’s adaptation of Bronfenbrenner’s Ecological Systems Theory (Murnane et al., 2018; Bronfenbrenner, 1974) highlights the importance of examining PHI within a network of health information ecologies, integrating formal health systems and informal social contexts. These frameworks have been extensively applied to domains like mental health (Tachtler et al., 2021), chronic conditions (Bhat and Kumar, 2020), PTSD/trauma (Evans et al., 2022; Wani et al., 2024), fertility (Costa Figueiredo and Chen, 2021), and substance abuse (Scroggs and Love, 2022), emphasizing how ecological approaches can inform technology design for managing physical, mental, and psychological aspects of an individual’s health and wellbeing.
Wearable technologies and smartphone applications are at the forefront of PHI in various contexts, like sports, managing chronic conditions, and broader long-term health monitoring initiatives (Tholander and Nylander, 2015; Klasnja et al., 2011; Poole, 2013; Nunes et al., 2015; Zegura et al., 2018; Blandford, 2019; Gooch et al., 2021; Fritz et al., 2014). Prior research on wearable use for PHI has also examined why people use and abandon smart devices (Lazar et al., 2015) and how it differs among people with prior self-tracking experience (Rapp and Cena, 2016). Studies on PHI often highlight the continuous capture and analysis of biosignals for precise intervention, particularly in chronic disease management, such as multiple sclerosis, chronic fatigue, and hypertension (Stuart et al., 2022; Jones et al., 2021; Tong et al., 2019).
While proactive use of PHI systems exists for critical use cases, such as emergencies (Amato et al., 2022; Alrowaily, 2024; Thulasi et al., 2024), fall detections(Delahoz and Labrador, 2014; Gjoreski et al., 2016; Pandya et al., 2020; Bastico et al., 2022; Kim et al., 2022), and medical reminders (Gjoreski et al., 2016; Brun and Häkkilä, 2022; Ma and Ren, 2023; Haque et al., 2024), we are yet to see them used for a broader PHI in the Indian context.
PHI in HCI is not limited to wearables and trackers; it also uses other modalities to integrate into an individual’s wellbeing. These diverse modalities represent different engagement strategies, from gamification in AR for improving physical activity (Zhao et al., 2024) using virtual pets to fostering social interaction through community displays (Gooch et al., 2021). Beyond physical and social wellbeing, wearables increasingly support mental health interventions through personalized, context-aware features (Ning et al., 2023; Kruzan et al., 2023). Despite this, the literature (Lee et al., 2022) often emphasizes specialized healthcare needs, such as technology readiness and e-health literacy, while general proactive health tracking remains underexplored in HCI, particularly in India. Our work adopts an ecological perspective to examine intrinsic and extrinsic factors influencing PHI adoption, bridging the gap between research on wearables and broader proactive health management strategies.
2.2. Barriers and Enablers of PHI in the Indian Context
@@
Kabir and Wiese’s meta-synthesis provides a broader understanding of the general global barriers and facilitators for personal informatics systems (Kabir and Wiese, 2023). Through this work, we explicitly explore India’s unique sociocultural and infrastructural challenges. Limited digital and health literacy among marginalized populations poses significant challenges. Older adults often face steep learning curves when adopting digital health technologies (Cherian et al., 2024), while children with ADHD struggle to interpret data from wearable devices, limiting meaningful engagement (Ankrah et al., 2023). Although simplified interfaces and accessible visualization techniques have been proposed to improve user engagement (Thakkar et al., 2022; Okolo et al., 2024; Lee et al., 2024b; Zhan et al., 2024; Wang et al., 2019; Mackert et al., 2016), additional supporting tools remain necessary to bridge the information gap.
@@
Efforts to improve engagement through public health information datafication have shown promise in enhancing accessibility and comprehension (Huang et al., 2022; Pai and Alathur, 2020). However, implementing these strategies in India demands policy-level changes to ensure equitable access and integration with local health systems. National healthcare initiatives—such as Ayushman Bharat Health Account (ABHA) and expanding telemedicine services—have gradually increased accessibility and awareness of PHI in India (Hwang et al., 2024; Napolitano et al., 2023). Engagement strategies through governmental schemes and NGO organizations, such as interactive, chat-based health information services and augmented displays, have shown promise in supporting medical professionals and older adults in managing their health (Karusala et al., 2023b; Sun et al., 2023; Luz and Masoodian, 2024). However, sustaining long-term engagement remains challenging, as users frequently cite usability issues and information overload as key reasons for abandonment (Lazar et al., 2015; Zhang and Shahriar, 2020; Lu et al., 2024).
@@
Apart from the Individual and Public/Infrastructural challenges, cultural and economic factors also play a pivotal role in shaping PHI adoption in India (Chandrasekaran et al., 2020; Upadhyay et al., 2023). Collectivist norms and uncertainty avoidance play a significant role in shaping attitudes toward digital health technologies (Karnatak et al., 2023; Zhang et al., 2022). The perceived economic value of health technologies is often underappreciated unless they provide critical utility to users through financial implications, for example, in managing chronic conditions (Singh et al., 2022b). While trust and perception issues persist, privacy concerns, data accuracy doubts, and discomfort with wearables further inhibit adoption (Shajari et al., 2023; Seelam et al., 2024).
@@
While these multi-faceted barriers exist in the Indian context, our study aims to reconfirm these, identify more, and develop strategies to improve health literacy, enhance PHI understanding and adoption, and sustain engagement by examining user trust, cultural awareness, and design personalization. Understanding and addressing these challenges will contribute to a more inclusive and sustainable adoption of PHI technologies.
@@ 3. Methodology
To examine the factors influencing the adoption of PHI in India, we conducted a study starting in January 2024, following approval from the Institute’s IRB. Our investigation began with a thorough review of the existing literature to understand the PHI usage and adoption landscape. It focused on the barriers and drivers for its adoption in personal health management and how its use (or the lack of it) is integrated into health management practices. We then designed a mixed-methods approach comprising surveys, interviews, and co-design workshops to capture the diverse perspectives of users. After analyzing all participant data and their inputs, we evaluated a rapid figma prototype of sharing preferences to further reinforce a unique finding.

### v2

# Unpacking Personal(?!) Health Informatics for Proactive Collective Care in India

@@ Abstract.
Personal Health Informatics (PHI), which leverages digital tools and information systems to support health assessment and self-care, promises more proactive, user-centered care, yet adoption and meaningful utilization barriers in India remain underexplored. Through a sequential mixed-methods study in urban India (Initial survey (n=87) and semi-structured interviews (n=22), follow-up survey = 116, and co-design workshops (n=6)), we surface practices, perceptions, and behaviors to identify ways PHI can be better utilized for proactive care in the Indian context. We find that PHI is valued for monitoring and enabling collective care; however, its adoption is constrained by low health and technology literacy, usability and integration issues, fragmented and costly technology ecosystems, and mistrust of digital health platforms. From triangulated evidence, we derive concrete design requirements, including user-controlled sharing, accessible analytics, and verifiable health information, and present a culturally grounded design vision for an integrated platform for collective care through design and evaluation of a figma prototype. The prototype evaluation provides further directions for design and development to better orient PHI for proactive care through the PHI-Proact operational map, which involves agency, elicitation, and engagement. Finally, using PHI-Proact, we conclude with concrete recommendations for designing and responsibly deploying PHI systems for proactive collective care in emerging contexts, which differ socially, culturally, infrastructurally, and technologically from WEIRD contexts.
@@ 1. Introduction
Personal health informatics (PHI) involves using information systems and tools to assess health and wellbeing and holds significant promise for empowering individuals and transforming healthcare by enabling proactive self-management. Globally, PHI leverages technologies (from smartphone applications to wearable devices such as smartwatches, fitness trackers, and other sensors), often facilitating continuous health monitoring (Shajari et al., 2023). Factors such as high health literacy, familiarity with technology, and a proactive healthcare ethos drive PHI adoption in Western contexts (Kyytsönen et al., 2023; Chandrasekaran et al., 2020; Zhang and Shahriar, 2020; Huang et al., 2022; Zhan et al., 2024; Zaher et al., 2024). Emerging evidence (Chmielewski and Meyer, 2024) in the Western context further suggests that healthcare providers and users are beginning to view PHI as a medium for a more engaged care, seeking personal insights and opportunities to share data with healthcare providers for more proactive management.
@@
In India, however, infrastructural, sociocultural, and technological challenges intersect and further constrain PHI’s potential for proactive care. Digital health tools, including wearables, are often perceived as fashion accessories or status symbols rather than integrated components of comprehensive health management (Misra et al., 2023). Moreover, cultural practices, family dynamics, and skepticism towards advanced technologies, such as AI-driven solutions, further influence PHI usage and adoption (Hoque and Bao, 2015; Zhang et al., 2022; Creaser et al., 2022; Patil et al., 2022; Reis et al., 2024; Kapania et al., 2022; Wisetsri and Vijai, 2021). Given the cultural dynamics of the Indian context, a significant gap in the adoption and use of PHI can be filled by examining and understanding how it is currently integrated and utilized in daily care practices in India and shared with family and loved ones.
@@
Guided by Murnane et al.’s ecological model of PHI (Murnane et al., 2018) and Bronfenbrenner’s Ecological Systems Theory (Bronfenbrenner, 1974), we investigate both intrinsic and extrinsic factors influencing PHI adoption in India for general wellbeing. Specifically, we ask (RQ1)How (if) do Indians adopt, integrate, and perceive PHI for proactive care? Based on the analyses of the initial components of the study, we further ask (RQ2) How can technologies be better aligned with lived practices to support proactive collective care?
To answer these questions, we investigated the adoption, integration, and perception of PHI among Indian users and stakeholders using a multi-method approach. We conducted a survey (n = 87) to assess awareness and usage of smart wearables, followed by interviews (n = 22) that explored user motivations, usage patterns, and sources of health information. After analyzing the collected data, our findings indicated a lack of awareness and access to tools that support their lived realities. To identify means to support that gap, we conducted additional interviews with key stakeholder pairs (couples, intergenerational family groups, and healthcare providers) and co-design workshops (n = 12) that employed card sorting, examining shared care practices and data-sharing preferences. Further, we evaluated a figma prototype to understand its usability and get design recommendations from the participants on what, how, and when they would like to share their health information/analytics with different stakeholders.
@@
Our findings provide a deeper understanding of PHI challenges and opportunities for proactive collective care in India, offering a Proof-of-Concept evaluation to further strengthen our recommendations for designers and developers. Through this work, we lay the groundwork for a more reliable and integrated PHI system that supports both personal and shared care practices and is culturally appropriate for non-WEIRD contexts, such as India.
@@ 2. Related Work
The body of work on PHI within the broader HCI community, though established, has seen significant expansion in recent years, driven by the rising prevalence of chronic diseases associated with sedentary lifestyles and advancements in digital health technologies (Anderson and Durstine, 2019). These factors have propelled research toward innovative health management solutions and broadened the adoption of PHI systems. The rapid increase in chronic health conditions and an expanding digital landscape make PHI especially critical in India. This work aims to contextualize the evolution of PHI towards collective care, addressing its unique challenges and opportunities. For the entirety of this work, we define collective care as the culturally grounded, trust-based system of shared responsibility for health and wellbeing that shapes information flow, decision-making, and support within a trusted circle of stakeholders (spouse, family, friends). This section examines the existing scholarship through two critical lenses: (1) the evolution of modeling and use cases of PHI in HCI, and (2) the barriers and enablers of PHI for Collective Care. These lenses offer complementary perspectives, highlighting technological and user-centric design considerations while incorporating them into India’s sociocultural care practices.
2.1. Evolution of Modeling and Use Cases of PHI in HCI
@@
Personal health informatics (PHI) models within HCI have evolved from individual-centric models (Rooksby et al., 2014; Zegura et al., 2018; Blandford, 2019; Kirchner et al., 2021) to stage-based models (Li et al., 2010) to broader models (Epstein et al., 2015; Pina et al., 2017) that account for social, structural, and cultural factors. While Murnane et al.’s adaptation of Bronfenbrenner’s Ecological Systems Theory (Murnane et al., 2018; Bronfenbrenner, 1974) highlights the importance of examining PHI within a network of health information ecologies, integrating formal health systems and informal social contexts for mental health, it also translates well to general health and wellbeing. These frameworks have been inherently or explicitly applied to domains like mental health (Tachtler et al., 2021), chronic conditions (Bhat and Kumar, 2020), PTSD/trauma (Evans et al., 2022; Wani et al., 2024), fertility (Costa Figueiredo and Chen, 2021), and substance abuse (Scroggs and Love, 2022) emphasizing how ecological approaches can inform technology design for managing physical, mental, and psychological aspects of an individual’s health and wellbeing. While these models support the individual even from an ecological perspective, there is a gap in understanding how all stakeholders can be supported to better utilize PHI for collective care contexts.
Moving on to applications, studies on PHI often highlight the usefulness of continuous capture and analysis of biosignals for precise intervention, particularly for specialized use cases. Earlier PHI systems developed using mobile phones and specialized sensors/technologies started with supporting older adults and people with movement disabilities during emergencies and fall detection (Ketabdar and Polzehl, 2009; Ko et al., 2010; Delahoz and Labrador, 2014; Robinson et al., 2015; Gjoreski et al., 2016; Pandya et al., 2020; Bastico et al., 2022; Kim et al., 2022; Amato et al., 2022; Alrowaily, 2024; Thulasi et al., 2024; Chang et al., 2024). However, over time, with advancements in wearable technology, its use has been expanded to provide medical reminders (Gjoreski et al., 2016; Brun and Häkkilä, 2022; Ma and Ren, 2023; Haque et al., 2024), activity monitoring/tracking/recognition (Shen et al., 2025; Chen et al., 2024; Le et al., 2024), and tracking behavior change (Intille, 2004; Klasnja et al., 2011; Sullivan and Lachman, 2017). Beyond these, PHI has evolved further to support chronic conditions (Raj et al., 2019; Tong et al., 2019; Stuart et al., 2022) monitoring continuous physiological signals to detect critical changes and also for other interventions with long-term impact goals such as fitness/self-improvment (Fritz et al., 2014; Lu et al., 2024; Wang et al., 2024), sports rehabilitation (Tholander and Nylander, 2015), and workplace wellness (Chung et al., 2017). These prior works provide a comprehensive understanding of how to model, comprehend, and utilize PHI, primarily in Western, Educated, Industrialized, Rich, and Democratic (WEIRD) contexts. However, a gap still exists in mapping the factors affecting its adoption in non-WEIRD collective care contexts.
While most of the examples given so far indicate the role of PHI through mobile and sensing-based approaches, PHI in HCI has evolved to support other modalities. These diverse modalities represent different engagement strategies: gamification in AR for improving physical activity (Zhao et al., 2024), utilizing virtual pets to foster social interaction through community displays (Gooch et al., 2021), and employing generative AI to provide personalized support tailored to the individual (Chopra et al., 2025). Beyond physical and social wellbeing, PHI (through wearables) increasingly supports mental health interventions through personalized, context-aware features (Ning et al., 2023; Kruzan et al., 2023; Jiang et al., 2023), thereby expanding the scope of PHI ever so slightly. Over the years, literature surveys (Nunes et al., 2015; Epstein et al., 2020; Motahar and Wiese, 2022) and case studies (Lin et al., 2025; Rey et al., 2023) have further added to this knowledge by keeping these critical factors and strategies for better utilization updated with the advances in technology. The work by Niess et al. (Niess et al., 2021) to examine beyond WEIRD user groups, and the work by Choube et al. on open-ended sensemaking of passive sensing (Choube et al., 2025) indicate a scope for tracking broader general wellbeing, there exists a gap in supporting various stakeholders in proactive collective care. Our work employs an ecological perspective to examine the intrinsic and extrinsic factors influencing PHI adoption, use, perception, and the stakeholders involved, with the aim of modeling it more effectively for proactive collective care. This approach bridges a gap in the research, particularly in non-Western, educated, industrialized, rich, and democratic (WEIRD) contexts, such as India.
2.2. Barriers and Enablers of PHI for Collective Care
@@
Kabir and Wiese’s meta-synthesis provides a broader understanding of the general global barriers and facilitators for personal informatics systems (Kabir and Wiese, 2023). Furthermore, despite advancements, there is an emphasis on technology readiness and e-health literacy for better adoption and utilization (Lee et al., 2022), as limited digital and health literacy, especially among marginalized populations, pose significant challenges. Older adults and specially-abled Individuals (such as blind or deaf individuals) often face steep learning curves, accessibility issues, privacy concerns, and more in their journey with personal health technologies (Zhang and Shahriar, 2020; Cherian et al., 2024; Lee et al., 2024c), and this barrier is amplified due to the lack of resources to support them in non-WEIRD contexts. Similarly, children with ADHD struggle to interpret data from wearable devices, limiting meaningful engagement (Ankrah et al., 2023). Although simplified interfaces and accessible visualization techniques have been proposed to improve user engagement and support self-compassion (Thakkar et al., 2022; Okolo et al., 2024; Loerakker et al., 2024; Lee et al., 2024d; Zhan et al., 2024; Wang et al., 2019; Mackert et al., 2016), additional supporting tools remain necessary to bridge the information gap where resources, knowledge, and infrastructure constraints amplify these barriers further in non-WEIRD contexts. Beyond these barriers, prior research on PHI use has further examined causes for disengagement (Park and Lee, 2023), bias in PHI systems (Yfantidou et al., 2023), adoption and usability (Lu et al., 2020), the lack of ”lifestyle fit” with individuals (Yang et al., 2015), why people use and abandon smart devices for self-tracking (Lazar et al., 2015; Jones et al., 2021), how it differs among people with prior self-tracking experience (Rapp and Cena, 2016), and other workarounds/customization for contextualization (Harrison et al., 2015; Feustel et al., 2018) to engage with PHI resourcefully. While understanding some of these barriers is resourceful for non-WEIRD contexts as well, they primarily support an individual and not the other stakeholders involved in collective care contexts. Through this work, we explicitly explore India’s unique sociocultural and infrastructural challenges towards supporting various actors involved in proactive collective care.
@@
Beyond the individual, broader efforts to improve engagement through the use of public health information datafication have also shown promise in enhancing accessibility and comprehension of health information in critical instances (Huang et al., 2022; Pai and Alathur, 2020). However, shared smartphone usage (as noted by prior studies (Gupta et al., 2022; Kaur and Singh, 2025; Aneja et al., 2025) among emergent Indian users) deters its application, and policy-level changes are required to ensure equitable access and integration with local health systems. National healthcare initiatives, such as Ayushman Bharat Health Account (ABHA) and the expansion of telemedicine services, have gradually increased accessibility and awareness of PHI in India (Hwang et al., 2024; Napolitano et al., 2023). Engagement strategies through governmental schemes and NGO organizations, such as interactive, chat-based health information services and augmented displays, have shown promise in supporting medical professionals and older adults, respectively, in managing their health (Karusala et al., 2023b; Sun et al., 2023; Luz and Masoodian, 2024). However, sustaining long-term engagement and implementing these strategies supporting multiple stakeholders in collective care remains challenging.
@@
Cultural and economic factors also play a pivotal role in shaping PHI adoption in India (Chandrasekaran et al., 2020; Upadhyay et al., 2023). Collectivist norms and uncertainty avoidance play a significant role in shaping attitudes toward digital health technologies (Karnatak et al., 2023; Zhang et al., 2022). The perceived economic value of health technologies is often underappreciated unless they provide critical utility to users through financial implications, for example, in managing chronic conditions (Singh et al., 2022b). While trust and perception issues persist, privacy concerns, doubts about data accuracy, and discomfort with wearables further inhibit adoption (Seelam et al., 2024).
@@
While these multifaceted barriers exist in the Indian context, our study aims to reconfirm these, identify additional ones, and develop strategies to improve health literacy/overcome the literacy gap, enhance PHI understanding and adoption, and sustain engagement by examining user trust, cultural awareness, and design personalization. Understanding and addressing these challenges will contribute to a more inclusive and sustainable adoption of PHI technologies for proactive collective care.
@@ 3. Methodology
We adopted a sequential exploratory mixed-methods design. We started the study in January 2024, following approval from the Institute’s IRB. RQ1 (survey + interviews) was exploratory and aimed to discover how urban Indian users understand, use, and leverage PHI for proactive care. Themes that emerged particularly around health sharing and collective care motivated a follow-up design phase. To investigate design responses to these emergent needs, we conducted co-design workshops, stakeholder interviews, prototype evaluation (RQ2), and a concluding survey. The co-design and prototype evaluation, therefore, functioned as a targeted, theory-informed design probe used to (a) generate design concepts grounded in participants’ existing practices, and (b) collect usability and preference feedback to refine those concepts. The workflow of the overall study is presented in Figure 1.